Mrs. Eubanks recieved $2000 as a wedding gift form her uncle. She wants to invest this money in a bank that will pay 4.5% interest compounded monthly. How much should she expect to have in the account after 20 years?
I am interested in knowing the formula to find the compounded interest and how to solve the answer.

Answer by stanbon(75887)

Mrs. Eubanks received $2000 as a wedding gift form her uncle. She wants to invest this money in a bank that will pay 4.5% interest compounded monthly. How much should she expect to have in the account after 20 years?
---------
A(t) = P(1+(r/n))^(nt)
--------
A(20) = 2000(1+(0.045/12))^(12*20)
----
A(20) = 2000(1+0.00375)^(240)
----
A(20) = 2000(2.4555)
----
A(20) = $4910.93
======================
Cheers,
Stan H.
======================